OpenStudy (anonymous):

please help... will fan and give medal solve the equation cos 2θ+3 sin 2θ=2, for 0◦<θ<90◦.

OpenStudy (anonymous):

expand cos2theta

OpenStudy (anonymous):

can you do the steps pls ??

OpenStudy (anonymous):

\[\cos2\theta = \cos ^{2}\theta - \sin ^{2}\theta\]

OpenStudy (anonymous):

can you do steps further to get the answer please ?

OpenStudy (cwrw238):

also use sin 2O = 2 sin O cos O

OpenStudy (anonymous):

yeah @cwrw238 plz help me

OpenStudy (cwrw238):

so we have cos^2 O - sin ^2 O + 6 sin O cos O - 2 = 0

OpenStudy (anonymous):

got it.. next ?

OpenStudy (cwrw238):

this is not as straight forward as i thought..

OpenStudy (anonymous):

yeah i know

OpenStudy (cwrw238):

i thought we might get a quadratic equation which we could easily solve but we cant

OpenStudy (anonymous):

yeah it wont end in a quadratid

OpenStudy (cwrw238):

one way would be to replace cos^2 x by 1 - sin^2 x and cos x by sqrt(1 - sin^ 2 x) which would be a bit messy....

OpenStudy (anonymous):

it will be too long i think

OpenStudy (anonymous):

@phi what do you think? could you help ?

OpenStudy (anonymous):

well you can do this problem by another way

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

consider a triangle with base of 3 units and height of 1 units

OpenStudy (anonymous):

the triangle will be a right angled triangle

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now the hypotenuse will be \[\sqrt{10}\]

OpenStudy (anonymous):

and the angle 71.6

OpenStudy (anonymous):